The element mercury, Hg, is a liquid at room temperature and has a density of 13.5 g/cm^3.

How many μg of mercury are in a 3.77 × 10^10 nL sample?

What is the volume, in mm^3 , of a 1.15 kg sample of Hg?
Answer #1

density =13.5 g/cm3

a) volume = 3.77 x 1010 nL

1nL = 10-6 cm3

therefore volume = 3,77 x 104 cm3

but density = mass/volume

mass = density x volume = 13.5 x 3.77 x 104 = 50.90 x 104 g

but 1 g = 106 ug

mass = 5.09 x 1010 ug

b)mass = 1.15 kg

but 1 kg = 1000 g

mass = 1150 g

density = mass/volume

volume = mass/density

volume = 1150 / 13.5 = 85.19 cm3

but 1 cm3 = 1000mm3

volume = 85.19 x 1000 = 85190 mm3